OpenStudy (anonymous):

Find the magnitude and direction angle of the vector v. v=8(cos135i + sin135j)

OpenStudy (anonymous):

That's magnitude \[8\sqrt{\sin ^2(135)+\cos ^2(135)}\] Remember trig identity!!